### v1

# How Does Perfect Fitting Affect Representation Learning? On the Training Dynamics of Representations in Deep Neural Networks

Abstract
In this paper, we elucidate how representations in deep neural networks (DNNs) evolve during training. We focus on overparameterized learning settings where the training continues much after the trained DNN starts to perfectly fit its training data. We examine the evolution of learned representations along the entire training process, including its perfect fitting regime, and with respect to the epoch-wise double descent phenomenon. We explore the representational similarity of DNN layers, each layer with respect to its own representations throughout the training process. For this, we use two similarity metrics: (1) The centered kernel alignment (CKA) similarity; (2) Similarity of decision regions of linear classifier probes that we train for the DNN layers. Our extensive experiments discover training dynamics patterns that can emerge in layers depending on the relative layer-depth, DNN width, and architecture. We show that representations at the deeper layers evolve much more in the training when an epoch-wise double descent occurs. For Vision Transformer, we show that the perfect fitting threshold creates a transition in the evolution of representations across all the encoder blocks.
1 Introduction
Deep neural networks (DNNs) have achieved remarkable practical success in various applications. The ability to learn representations for each of the DNN layers is an essential aspect that determines the overall DNN performance. Yet, the inner workings of the representation learning process are still far from being sufficiently understood. An important question that we address in this paper is how do the training dynamics of the representations behave in each of the DNN layers?
Usually, DNNs are overparameterized models, namely, they have many more learnable parameters than training examples. Due to such overparameterization, DNNs are often trained to perfectly fit their training data (e.g., correctly labeling all the training examples in a classification task). This perfect fitting property stands in sharp contrast to conventional machine learning guidelines that associate overfitting (and, hence, perfect fitting) with poor generalization performance. Yet, DNNs that perfectly fit their training data often show good generalization performance, raising foundational questions on their inner workings.
The training process of overparameterized models can be divided into the following two parts. The first part is from the beginning of training (initialization) and until the time point (training epoch) when the model perfectly fits its training data for the first time. Then, the second part is the perfect fitting regime of training where the trained model always perfectly fits its training data (Fig. 1g). Evaluating the test error of the model as a function of the training epoch can show (especially when label noise exists) an epoch-wise double descent phenomenon [1] that includes two parts (Figs. 2g,7g): (i) A classical “U-shape” that reflects the classical approach of an ideal early stopping point: the test error first decreases, then increases back and peaks towards the start of the perfect fitting regime. (ii) A decreasing test error after its peak and into the entire perfect fitting regime that can eventually achieve a low test error, suggesting the approach of long training as an alternative to early stopping.
In this paper, we ask how do overparameterization and perfect fitting affect the evolution of representations in each of the DNN layers during training? To address this question, we evaluate how similar are the representations that evolve at each layer along the entire training process. Namely, for a layer of interest, we provide a detailed map that shows how similar are the representations of that layer at every two different time points (epochs) in the training. We believe that our evaluations provide the most extensive and detailed elucidation of representational similarity in the training process (as we discuss in Section 2, some previous works had a very limited evaluation of representational similarity in training with respect to only a particular training time-point).
In our evaluations, we use two different approaches to quantify representational similarity. First, the centered kernel alignment (CKA) [2], which is a popular measure of representational similarity between layers of different DNNs or within the same DNN. Second, we propose a new approach to measure representational similarity by using linear classifier probes [3] for the DNN layers during their training process. We use the linear classifier probes of the same layer at different training epochs and evaluate the similarity of their decision regions as a measure of similarity between the representations at these training epochs.
Using our detailed representational similarity evaluation, we examine overparameterized learning settings where the training continues much after the DNN starts to perfectly fit its training data. By this, we elucidate the evolution of representations at the various parts of training, including the perfect fitting regime, and in settings with/without the epoch-wise double descent phenomenon. This sheds new light on representation learning in overparameterized models, including the following insights:
  1. 1.
    Representations at the first layer, in the perfect fitting regime of training a sufficiently wide DNN, are more similar to random representations (e.g., as at the initialized, untrained DNN) than to representations at the best early stopping epochs before perfect fitting starts.
  2. 2.
    Representations at the deeper layers evolve much more in the training when an epoch-wise double descent occurs. This is especially apparent in ResNet training.
  3. 3.
    Only the linear classifier probes of the deeper layers have an epoch-wise double descent (in their own, probing test error) when the entire DNN also has the epoch-wise double descent. This suggests that the representations and features at the deeper layers have a crucial role in creating the epoch-wise double descent phenomenon of the entire DNN.
  4. 4.
    Representations often have one or two transitions in their training evolution. These create training segments in which representations are relatively similar, visually appearing as lighter color “similarity blocks” along the main diagonal of the similarity diagrams. For ResNet, this is especially visible when label noise and epoch-wise double descent exist (see, e.g., Figs. 2c-2f). For Vision Transformer (ViT), this is well observed for both the noiseless and label noise settings (Figs. 3, 17); moreover, we show that the perfect fitting threshold creates a transition in the evolution of representations across all the ViT encoder blocks.
2 Related Work
The effect of DNN width on similarity of representations was examined in [4] using the CCA metric, and in [2, 5] using the CKA metric. In contrast to these works, our focus is on elucidating the representional similarity during the training process. In the following, we discuss the existing literature which is more related to our work.
2.1 Representational Similarity Throughout the Training Process
In [6], the SVCCA metric for representational similarity is proposed and utilized for several analyses, including showing how the representational similarity between the model during training and the model at the end of training evolves during training of a convolutional model and a ResNet model. In addition, [7] shows the evolution of the SVCCA metric during 4 epochs of a BERT model training. In this paper, we examine the evolution of representational similarity much more extensively than previous works, we consider various new settings (including for a ViT architecture, noiseless and noisy datasets) and, most importantly, we use it for the first time as a tool for analyzing the perfect fitting effect and the epoch-wise double descent phenomenon.
In [8], the loss landscape (specifically, connectivity of local minima) and representational similarity (using the CCA metric [6]) are examined for various training heuristics from the common practice, such as learning rate restarts, warmup, and distillation. Their results include evaluation of CCA representational similarity between a VGG model at training iteration 0 and the specific training iteration at the end of their warmup phase (if applied); their evaluation reflects high similarity for the first convolutional layer of the model. They also evaluate the CCA similarity between the model at the end of training and at the intermediate training iteration after the warmup. Importantly, the work in [8] does not show or analyze the detailed evolution of the representational similarity during the entire training process. In contrast, in our paper here we examine the representational similarity in much more detail, enabling to understand the representational similarity of various pairs of epochs, throughout a long training process with a small constant learning rate that does not affect the natural training dynamics. Moreover, we focus here on understanding how the representational similarity evolves in overparameterized learning where perfect fitting does not prevent good generalization.
2.2 Representational Similarity between Trained and Untrained Models
In [2], the CKA similarity between trained and untrained convolutional DNNs was examined. They show that the representations in the first layers of the DNN are similar to those of the corresponding layers in the untrained DNN. Their CKA values for the first layer of a 10-layer convolutional model on CIFAR-10 was slightly above 0.8. In this paper (e.g., Figs. 1a, 2a, 7a), we show that the first layer can have CKA similarity (nearly) 1 in the perfect fitting regime of training.
In Appendix E we discuss the related work [9] and explain how we disprove the generality of their claim regarding the learning process of the first layer in the DNN. We also discuss in Appendix E an additional related work [10].
Importantly, in addition to our extensive CKA evaluations, we propose a new representational similarity measure based on similarity of the decision regions of linear classifier probes. Both the CKA and the new probing-based measure lead to new insights on representation learning that are not available in the existing literature.
3 Training Dynamics of Representational Similarity using CKA
3.1 The CKA Measure: Definition and Formulation
Quantifying similarity of representations in intermediate DNN layers is known as a challenging task due to the inconsistent and intricate role the various neurons have in making the representations. The CKA was proposed in [2] as a robust method for measuring similarity of representations; indeed, the CKA (based on a linear kernel) is a popular, leading choice for analyzing representations in DNNs (e.g., [5, 11]).
Consider two classification DNNs, denoted as the functions $f,g:\mathcal{X}\rightarrow\mathcal{Y}$ where $\mathcal{X}$ is the input domain and $\mathcal{Y}$ is the output domain (set of possible labels). We denote the mapping from the input to the representation of the $L^{\sf th}$ layer (i.e., the activations of the $L^{\sf th}$ layer) of a DNN $f$ as $f_{L}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$. Here, $p_{L}$ is the number of neurons in layer $L$ and, hence, also the dimension of the layer’s representation when organized as a column vector.
The CKA is computed for two layers of possibly different DNNs, namely, $f_{L}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$ and $g_{L^{\prime}}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L^{\prime}}}$ based on a set of $m$ inputs examples $\mathcal{S}_{m}=\{\mathbf{x}_{1},...,\mathbf{x}_{m}\}\subset\mathcal{X}$. For these examples, the representations are computed and organized in the following matrices: $\mathbf{F}\triangleq\left[f_{L}\left(\mathbf{x}_{1}\right),...,f_{L}\left(\mathbf{x}_{m}\right)\right]^{T}$ and $\mathbf{G}\triangleq\left[g_{L^{\prime}}\left(\mathbf{x}_{1}\right),...,g_{L^{\prime}}\left(\mathbf{x}_{m}\right)\right]^{T}$. Then, $\mathbf{K}\triangleq\mathbf{F}\mathbf{F}^{T}$ and $\mathbf{L}\triangleq\mathbf{G}\mathbf{G}^{T}$ are the Gram matrices of the representations (separately for $f_{L}$, $g_{L^{\prime}}$) whose components quantify the representation similarity between pairs of examples. These Gram matrices further go through centering of their rows and columns, namely, computing $\mathbf{K}^{\prime}=\mathbf{H}\mathbf{K}\mathbf{H}$ and $\mathbf{L}^{\prime}=\mathbf{H}\mathbf{L}\mathbf{H}$ where $\mathbf{H}\triangleq\mathbf{I}_{m\times m}-\frac{1}{m}\mathbf{1}_{m\times m}$ is the centering matrix, $\mathbf{I}_{m\times m}$ is the $m\times m$ identity matrix, and $\mathbf{1}_{m\times m}$ is an $m\times m$ matrix of ones. Then, the HSIC is computed as the normalized dot product of the vectorized forms of $\mathbf{K}^{\prime}$, $\mathbf{L}^{\prime}$:
${\sf HSIC}_{0}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})=\frac{1}{m-1}{\sf vec}\left(\mathbf{K}^{\prime}\right)\cdot{\sf vec}\left(\mathbf{L}^{\prime}\right)$ (1)
where $\mathbf{K}^{\prime}$ and $\mathbf{L}^{\prime}$ were defined above based on $f_{L}$ and $g_{L^{\prime}}$, respectively, and the set of examples $\mathcal{S}_{m}$. Importantly, the HSIC measure is not affected by permutation of neurons and orthogonal transformations of the representations. The important property of invariance to uniform scaling of the representations is added by the following normalization that defines the CKA measure:
${\sf CKA}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})=\frac{{\sf HSIC}_{0}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})}{\sqrt{{\sf HSIC}_{0}(f_{L},f_{L};\mathcal{S}_{m})}\sqrt{{\sf HSIC}_{0}(g_{L^{\prime}},g_{L^{\prime}};\mathcal{S}_{m})}}.$ (2)
3.2 The Proposed CKA Evaluation During the Training Process
Figure 1: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 without label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.
Figure 2: CKA evaluations for ResNet-18 (standard width), trained on CIFAR-10 with 20% label noise.
In this section we examine the evolving representations of the DNN layers throughout the training. The similarity of representations in each layer are computed with respect to the representations of the same layer, at different time points (epochs) in the training process. This can be formulated as follows. The mapping from the input to the representation of the $L^{\sf th}$ layer at training epoch $t\in\{0,\dots,3999\}$ is denoted as $f_{L}^{(t)}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$, where $p_{L}$ is the dimension of the representation of layer $L$. Then, we compute the CKA for $f_{L}^{(t)}$ and $f_{L}^{(t^{\prime})}$ for every two training epochs $t,t^{\prime}\in\{0,\dots,3999\}$ [^1] In experiments on ResNet and CIFAR-10, we always evaluated the entire detailed epoch grid from epoch 0 to 3999. In the other experiments, we downsampled the epoch grid for which the CKA is computed, due to the high computational and storage demands; even then, our epoch grid is relatively dense w.r.t. the presentation of the results using a logarithmic epoch axis. The details on the epoch grids are provided in Appendix F.. Namely, for a layer $L$, we create a 2D diagram (heatmap) of CKA values, ${\sf CKA}(f_{L}^{(t)},f_{L}^{(t^{\prime})};\mathcal{S}_{m})$ for $t,t^{\prime}\in\{0,\dots,3999\}$.
We trained ResNet-18 and ViT-B/16 DNNs on the CIFAR-10 dataset (without and with 20% label noise). The ResNet-18 was trained also on the Tiny ImageNet dataset. Unless otherwise specified, models were trained for 4000 epochs with the Adam optimizer at a constant learning rate of 0.0001 (without weight decay) that ensures smooth and natural evolution of the training process (this was also the leading setting in the deep double descent study in [1]). In addition, we provide exemplary results for the SGD optimizer with/without momentum (Appendix Figs. 13, 14, 15), and for Adam with weight decay (see Appendix Fig. 16). Standard data augmentation was applied. Random initialization was applied, epoch 0 corresponds to the initialized DNN before its training starts. See Appendix G for the summary of experiment settings.
The resulting CKA diagrams are provided in Figs. 1, 2, 3 and in Appendix Figs. 7-17. Note that the CKA of two training epochs is symmetric, ${\sf CKA}(f_{L}^{(t)},f_{L}^{(t^{\prime})};\mathcal{S}_{m})={\sf CKA}(f_{L}^{(t^{\prime})},f_{L}^{(t)};\mathcal{S}_{m})$. Naturally, the similarity is maximal along the main diagonal that connects the lower-left and upper-right corners, as values along this diagonal correspond to similarity of a layer at a specific epoch with itself. Usually, the similarity is relatively high around the main diagonal, as representations usually evolve relatively smooth throughout the training process.
Each of the 2D diagrams considers a specific layer in the DNN and can be read according to its rows. Each row is associated with a different training epoch to consider its representation as a reference to evaluate the similarity for the entire training process, which corresponds to the horizontal axis of the diagram. For example, the first (bottom) row of the diagram shows how the representations throughout the entire training process are similar to the random representations at the initialization. In the following, we discuss the various findings that stem from these CKA diagrams.
3.3 Analysis of the CKA Results
Figure 3: CKA evaluations for ViT-B/16, trained on CIFAR-10 without label noise.
Representations at the first layer are more similar to random representations than to good early-stopping representations.
The CKA diagrams of the first layer in ResNet-18 trained on CIFAR-10 without/with label noise (Figs. 1a, 2a) show higher CKA values than the other layers, suggesting that representations of the first layer do not change much compared to their random initialization. Yet, within this range of high CKA values, the diagrams show that in the perfect fitting regime of training the first layer representation is more similar to the random representations at the initialization and early in training rather than to representations in the beneficial early-stopping range of training (i.e., at epochs where the test error is relatively low and perfect fitting did not start yet, according to Figs. 1g, 2g). Note that the similarity to the random representations is not only to the specific initialization of the training process, but also to other random representations unrelated to the specific training process; we examine this further in Appendix H.1 and Fig. 19. While most of our results are for the Adam optimizer, for the SGD optimizer (Figs. 13a, 14a, 15a) the first layer representations behave somewhat differently, although they still have very high CKA similarity during the entire training process.
Importantly, the representations of the first layer have increased similarity to the random initialization at the perfect fitting regime only if the DNN is wide enough. For example, see Figs. 8a, 9a where we examine ResNet-18 of smaller widths (width parameters $k=32,42$) that still achieve perfect fitting (and show epoch-wise double descent) but do not have increased similarity to initialization at the end of training. When the DNN is wider (width parameter $k=52$), in Fig. 10a, the behavior is more similar to our results in Fig. 2a for the standard width of ResNet-18 (width parameter $k=64$).
Training the standard-width ResNet-18 using weight decay regularization with the Adam optimizer (Fig. 16a) shows similar behavior of the first layer representations as for smaller width ResNet-18 trained without weight decay (e.g., Fig. 10a). This can be explained by the effective reduction of the model complexity due to the weight decay regularization, making it similar to the model complexity of unregularized smaller model. Note that the weight decay regularization lowers the CKA similarity of the first layer representations (Fig. 16a) in a few percentages compared to training without weight decay (even of smaller width ResNet-18 models); this suggests that weight decay can increase to some extent the evolution of the first layer representations.
Deeper layers evolve much more when epoch-wise double descent occurs.
Figures 1, 2 consider the same learning setting for ResNet-18, except for the single difference of having noiseless versus 20% label noise in the CIFAR-10 dataset. Accordingly, an epoch-wise double descent in the test error occurs only for the noisy case (compare Figs. 1g, 2g). Visually comparing the CKA diagrams of the two cases (Figs. 1,2) shows that the representations at the deeper layers evolve much more in the training when an epoch-wise double descent occurs. This is visually emphasized by the darker color (i.e., lower similarity) in the heatmaps of the fourth ResNet Block (Figs. 1e, 2e) and the last FC layer (Figs. 1f, 2f), reflecting lower CKA similarity among nearby training epochs due to excessive instability and/or evolution of the representations of these layers.
Representations can have one or two transition points in their training evolution, creating “similarity blocks” along the main diagonal of the similarity diagram.
For ResNet, this behavior is especially visible in settings with label noise and epoch-wise double descent (Fig. 2). For example, see Figs. 2c-2f where at each layer “similarity blocks” of relatively lighter color appear around the main diagonal of the diagram. Importantly note that different ResNet layers can differ in the number of “similarity blocks” as well as the range of epochs in which the “similarity blocks” emerge. Specifically, the earlier ResNet layers (Figs. 2b-2d) have larger “similarity blocks” that do not seem to be affected by the perfect fitting regime of training; in contrast, the deeper layers (Figs. 2e-2f) have smaller (or medium size) “similarity blocks” that seem to be more affected by the perfect fitting regime (in particular note that the last FC layer has a small “similarity block” towards the end of training which is very dissimilar to the entire training process; see this also for ResNet-50 in Fig. 12f). In the noiseless setting (where there is no epoch-wise double descent) of Fig. 1, the “similarity blocks” appear more weakly than in the noisy case; this emphasizes the more segmented nature of representation learning in ResNet when the epoch-wise double descent occurs. Yet, the in the noiseless case of ResNet-18, there are “similarity blocks” that are aligned with the perfect fitting regime (i.e., they start near to epoch 1000 where the perfect fitting starts, see Figs. 2e, 2f). Moreover, interestingly, training ResNet-18 with SGD (Figs. 13, 14, 15) has clearer “similarity blocks” and transition points between them than in training with Adam.
ViT has clear “similarity blocks” and transition points which are aligned with the perfect fitting regime.
The phenomenon of “similarity blocks” is somewhat visibly clearer in ViT than in ResNet. Specifically, in ViT training the “similarity blocks” are well observed for both noiseless (Fig. 3) and noisy (Fig. 17) datasets. Specifically, the ViT encoder blocks at various depths have “similarity blocks” which are aligned with the perfect fitting regime (i.e., they start around epoch 1000 where perfect fitting also starts; see Figs. 3e, 3f, 17j, 17m, noting that at the deep layers these “similarity blocks” are visibly clearer).
4 Decision Regions Similarity of Linear Classifier Probes of Layers During Training
Following the insights we obtained using the CKA evaluations, we now turn to a second, significantly different, approach to analyze the training dynamics of the learned representations and features in the DNN layers. Specifically, in this section, we propose a new analysis approach that can be of independent interest also beyond the research questions of this paper.
4.1 Training Linear Classifier Probes for Layers Throughout the Training Process
From the complete training process, we define a relatively dense subset of training epochs $\mathcal{T}\subset\{0,1,\dots,3999\}$ for which we examine various layers in the trained DNN. As the notations earlier in this paper, for layer $L$ at epoch $t\in\mathcal{T}$, the mapping from the input to the layer’s representation is $f_{L}^{(t)}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$. We use $f_{L}^{(t)}$ (i.e., the DNN segment from the input to layer $L$) as a fixed feature map, and add a trainable linear layer on top of it to act as a probe. Namely, we train the weights $\mathbf{W}\in\mathbb{R}^{|\mathcal{Y}|\times p_{L}}$ and bias $\mathbf{b}\in\mathbb{R}^{|\mathcal{Y}|}$ for the linear classifier probe
$\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf softmax}\left({\mathbf{W}}f_{L}^{(t)}\left(\mathbf{x}\right)+\mathbf{b}\right)$ (3)
where $\mathcal{Y}$ is the label domain of the original classification task. The training of the linear probe is completely separate from the training of the original DNN that $f_{L}^{(t)}$ is taken from. Moreover, linear probes at different training epochs (of the original DNN) are trained separately and do not share their weights $\mathbf{W}$ and bias $\mathbf{b}$. See more details on the probe training in Appendix F.3. Note that the linear probe layer in (4) maps the $p_{L}$-dimensional representation of $f_{L}^{(t)}$ to a score vector of classes (these classes are as in the classification task of the original DNN, and the probe’s score vector also goes through a softmax function). The predicted class label from $\mathcal{Y}$ is the one with the maximal softmax-based class score, namely, by denoting the $j^{\sf th}$ class score as $\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)[j]$ we have
${\sf Probe}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf argmax}_{j\in\{1,\dots,|\mathcal{Y}|\}}\widetilde{{\sf Probe}}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)[j].$ (4)
4.2 Measuring Similarity of Decision Regions of Linear Classifier Probes
Figure 4: Decision regions similarity of linear classifier probes. Evaluations for ResNet-18 (of standard width), trained on CIFAR-10 without label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The error curves are provided in Fig. 1g.
Figure 5: Decision regions similarity (DRS) of linear classifier probes. Evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The test and error curves are provided in Fig. 2g.
Linear classifier probes are often used for analyzing the quality of the learned features by evaluating their usefulness for the original [3, 11] or another classification task [10]. In this paper, we are interested in understanding the evolution of representations during training; accordingly, we propose here a new probing-based similarity measure that reflects the overall representation. Specifically, we propose to compare representations of the same layer at two different training epochs by evaluating the similarity of the decision regions of the linear classifier probes of the layer at the respective training epochs.
Evaluating similarity of decision regions was proposed in [12] for comparing fully-trained DNNs in their standard classification functionality of the entire DNN (i.e., not its layers and without probes). Here, we extend this decision-region similarity evaluation to the completely different analysis task of comparing linear probes that correspond to intermediate layers during their training.
Given two linear classifier probes of the same layer $L$ at different training epochs $t,t^{\prime}$, we numerically estimate the relative portion of the input space for which their predictions agree. As in [12], we define $N_{Q}=500$ planes that reside in the input space, each plane is defined by randomly choosing three training examples and computing the plane that they span. See more details in Appendix F.3. The $j^{\sf th}$ plane, $j\in\{1,\dots,N_{Q}\}$, is denoted as $\mathcal{Q}_{j}\subset\mathcal{X}$. $\mathcal{Q}_{j}^{\sf discrete}$ is a discrete, finite version of $\mathcal{Q}_{j}$, which is defined for the numerical evaluation based on a uniformly-spaced 2D grid of $|\mathcal{Q}_{j}^{\sf discrete}|=2500$ points. Each $\mathbf{x}\in\mathcal{Q}_{j}^{\sf discrete}$ is an image, sampled from the plane, that can be given to $f_{L}^{(t)}$ and $f_{L}^{(t^{\prime})}$ as input. Then, the similarity of the decision regions of the probes of layer $L$ at epochs $t,t^{\prime}$ is
${\sf DRS}\left(f_{L}^{(t)},f_{L}^{(t^{\prime})}\right)\triangleq\frac{1}{N_{Q}}\sum_{j=1}^{N_{Q}}\sum_{\mathbf{x}\in\mathcal{Q}_{j}^{\sf discrete}}\frac{\mathbb{I}\left({\sf Probe}_{L}^{(t)}\left({f_{L}^{(t)}\left(\mathbf{x}\right)}\right)={\sf Probe}_{L}^{(t^{\prime})}\left({f_{L}^{(t^{\prime})}\left(\mathbf{x}\right)}\right)\right)}{|\mathcal{Q}_{j}^{\sf discrete}|}$ (5)
where $\mathbb{I}({\sf condition})$ is an indicator function that returns 1 if the ${\sf condition}$ is satisfied, and 0 otherwise. According to (5), the value of ${\sf DRS}\left(f_{L}^{(t)},f_{L}^{(t^{\prime})}\right)$ is between 0 and 1, where a higher value reflects higher similarity and agreement of the decision regions of the linear probes.
The foundational idea that we propose here is that similar representations lead to similar decision regions of the respective linear classifier probes. In this approach, the intricacy of comparing the representations of different models (from different training epochs) is first addressed by the training of linear classifier probes, which interpret the representational structure to achieve their own training objectives. Clearly, the functionality that the probes learn reflect and depend on the representational structure of their inputs, while having the good merits (for representational similarity evaluation) of insensitivity to the particular order of neurons, orthogonal transformations, or uniform scaling (it should be noted though that there is some level of inherent randomness in the training process of the probes). Comparing the decision regions of the probes is a way to compare their functionality and, by that, to compare the representational structure they get as their inputs. Clearly, this is an inherently different approach to measure representational similarity than using CKA as in Section 3; indeed, we will now see it can shed additional light on the evolution of representations in training and further support insights we obtained using the CKA.
4.3 Analysis of the Linear Classifier Probes Results
In Figs. 4, 5, 18 we show results of the decision regions similarity (DRS) of the linear classifier probes, each diagram corresponds to a different layer in the DNN. For a start, note that the DRS evaluations in Figs. 4, 5 support insights that we got from the CKA in Figs. 1, 2. Specifically, note in Figs. 4a, 5a that towards and in the perfect fitting regime of training, the first layer becomes more similar to random representations as it had at the initialization. Yet, there are two main differences here compared to the CKA results. First, the DRS values have a wider range than in the CKA for the first layer, suggesting that indeed there is a significant training dynamic where the first layer representations become more similar to random representations. Second, the DRS shows that the similarity to random representations towards and in the perfect fitting regime is temporary and eventually at the end of training the first layer representation is different than for random representations. The ViT results in Fig. 18 show a different behavior than ResNet, but this was also the case for the CKA evaluations.
Moreover, comparing the DRS evaluations in Figs. 4d, 4e to Figs. 5d, 5e supports our CKA-based insights that when an epoch-wise double descent occurs, the representations at the deeper layers in the DNN are more actively evolving (i.e., the DRS values at the deeper layers are lower when epoch-wise double descent occurs).
Figure 6: Test error of the linear classifier probes during a ResNet-18 training on CIFAR-10 (a) without label noise, (b) with 20% label noise. Each curve color refer to another layer in the DNN.
Importantly, our linear classifier probes here show that an epoch-wise double descent is reflected only by the linear classifier probes at the deeper layers. See the test error curves of the linear classifier probes as function of the training epoch of the original DNN (Fig. 6). Specifically, in the experiment with label noise (and epoch-wise double descent), only the probe of the fourth ResNet block shows a significant epoch-wise double descent (Fig. 6b). This suggests that the epoch-wise double descent phenomenon crucially depends on the deeper layers and their fine grained features.
5 Conclusion
In this paper we have elucidated the training dynamics of representations in deep learning. We have used the CKA similarity measure and proposed a new similarity measure based on decision regions of linear classifier probes. Our extensive experimental evaluations provide a highly detailed description of the representation evolution during training. By this, we provide several new insights on the representation learning dynamics, with a particular focus on overparameterized learning that continues much after perfect fitting starts. We believe that the new insights and empirical analysis tools of this paper can pave the way for additional works on related research questions.
Appendices
Appendix A Limitations
In this paper we study the representational similarity during training of DNNs. We use the CKA similarity measure and propose a new similarity measure based on the decision regions of linear classifier probes. Measuring representational similarity is known as an intricate task, and indeed any similarity measure has its own strengths and weaknesses. Accordingly, while our results provide new insights, most of which are confirmed by the two similarity measures we use, other similarity measures may be proposed in the future and lead to additional insights and discoveries.
Moreover, we present in this paper a highly extensive, computationally-intensive set of experiments on various training settings. Nevertheless, we focus on the overparameterized learning settings where the trained model is able to perfectly fit its training data, and the training continues much after the point where perfect fitting starts; our focus on these learning settings conforms with the research questions that we address in this paper and related to the effect of overparameterization and perfect fitting on the representation learning process. While we provide an exemplary experiment for training with explicit regularization (Adam with weight decay), future work can consider in more detail regularized training settings and search for new insights for them.
Appendix B Broader Impact
This paper presents a foundational research of the training process of DNNs. As such, the understanding and insights in this paper may reduce to some extent the extensive trial and error engineering in the common practice and, by that, may reduce the amount of time and energy needed for developing new deep learning models. Moreover, foundational insights, as in this paper, can potentially contribute to better interpretability of deep learning models.
Appendix C Experiments Compute Resources
For all of our experiments, we utilized an internal cluster equipped with Nvidia RTX 3090 and RTX 4090 GPUs, each with 24GB of memory. Below is a summary of the compute resources and run times for each experiment:
  • Model training:
    • ResNet-18 and ViT-B/16 models trained on the CIFAR-10 dataset: Approximately 15 hours.
    • ResNet-18 and ViT-B/16 models trained on the Tiny ImageNet dataset: Approximately 48 hours.
  • CKA similarity heatmap diagrams:
    • Generation of CKA similarity heatmap diagrams: Approximately 15 hours per model, using 6 GPUs.
  • Decision region similarity (DRS) of linear classifier probes:
    • Generation of DRS heatmaps: Approximately 3 days per model, using 12 GPUs.
In terms of the number of experiments:
  • Training base models:
    • We trained approximately 10 models of ResNet-18 on CIFAR-10:
        4 different widths (32, 42, 52, 64).
        3 different label noise levels (with width parameter k=64).
        2 models were trained with SGD, with and without label noise (all the other models were trained with the Adam optimizer).
        1 model was trained with weight decay regularization.
    • We trained 2 ResNet-18 models on the Tiny ImageNet dataset, with and without label noise.
    • We trained 1 ResNet-50 models on the CIFAR-10 dataset, with 20% label noise.
    • We trained 2 models of ViT-B/16 on the CIFAR-10 dataset, with and without label noise.
  • CKA similarity heatmap diagrams:
    • For each model we trained, we also produced the CKA heatmap diagram, resulting in a total of approximately 14 CKA heatmap experiments.
  • Decision region similarity (DRS) of linear classifier probes:
    • We produced DRS heatmaps for ResNet-18 with 20% label noise, without label noise, and for ViT-B/16 without label noise, resulting in a total of 3 DRS heatmap experiments.
In terms of storage requirements:
  • For storing ResNet-18 models, we stored a checkpoint of the model after every epoch of the 4000 epochs, requiring approximately 170GB of storage per experiment (for all 4000 epochs).
  • For storing ViT-B/16 models, we stored a checkpoint of the model following this grid: all epochs from 0 to 300, every third epoch between 300 to 900, and every tenth epoch from 900 to 3999, resulting in a total of 350GB per experiment.
  • In total, we used around 4.5 TB of storage for all experiments described in this paper.
We did not use significant compute resources for any other tasks beyond the specified above.
Appendix D Assets and Software Packages
For the datasets, we used CIFAR-10 and Tiny ImageNet, which are publicly available:
  • CIFAR-10: Available at https://www.cs.toronto.edu/~kriz/cifar.html. The CIFAR-10 dataset includes 10 classes, each with 5000 training examples, and a total of 50,000 training examples.
  • Tiny ImageNet: A subset of 200 classes out of the ImageNet dataset. 500 training examples per class, and a total of 100,000 training examples. We used a downsampled version of $32\times 32\times 3$ pixels as inputs to our models.
For software, we utilized several packages:
  • PyTorch: Used for training the DNNs and running computations on GPUs. PyTorch is distributed under the Apache Contributor License Agreement (CLA). More information can be found at https://pytorch.org/.
  • torch-cka: Used for CKA computation, incorporating technical methods from the package. The package is available at https://github.com/AntixK/PyTorch-Model-Compare and is licensed under the MIT License.
  • dbViz: For decision region similarity (DRS) experiments, we used methods introduced in [12] with code available at https://github.com/somepago/dbViz, licensed under the Apache 2.0 License.
Appendix E Additional Related Work
This Appendix extends the Related Work section of the main paper (Section 2).
In [9], the authors propose to analyze the first layer in convolutional DNNs using the energy spectrum of the learned filters in the PCA basis of the input image patches. Their results show a typical form of the learned energy spectrum which is consistent over different convolutional architectures and datasets but very different compared to its random initialization at the beginning of training; we disprove the generality of this claim in this paper, by using the CKA similarity metric and showing that the representations of the first layer of a ResNet model can be very similar to the representations at the random initialization. This difference is probably due to our training setting that allows long training in the perfect fitting regime and avoids early stopping (whereas their training is much shorter; for example, their Fig. 18c in [9] shows a training process of 200 epochs that stops approximately when perfect fitting starts).
In [10], the authors use synthetic datasets (with controllable feature relevance) to study which features a trained model uses and how the choice of the features relates to their relevance to the task and to their linear-decodability from representations in an untrained model. Moreover, they show that trained models have relatively higher representational similarity (quantified by the neuroscientific method from [13]) to untrained models than to models trained on another task. Importantly, the representational similarity between trained and untrained models which is reported in [10] is moderate, and even quite low, in contrast to our results in this paper that shows much higher similarity levels according to other similarity metrics.
Appendix F Additional Details on the Experiment Settings
F.1 Models
ResNet-18: We used the (Preactivation) ResNet-18 architecture as introduced in [14]. The ResNet-18 structure includes an initial convolutional layer followed by four ResNet blocks, each consisting of two BatchNorm-ReLU-Convolution layers. The layer widths for these four ResNet blocks are defined as [k, 2k, 4k, 8k] where $k$ is a width parameter (positive integer), and $k=64$ corresponds to the standard ResNet-18 layout. Similarly to the deep double descent research in [1], we consider several values of $k$ to examine the effect of width and parameterization level on the ResNet-18 model. The ResNet-18 implementation was as in [1, 12] that originate in https://github.com/kuangliu/pytorch-cifar.
ResNet-50: For the experiment on ResNet-50 in Fig. 12 we used a standard ResNet-50 implementation that originates in https://github.com/kuangliu/pytorch-cifar and follows [14]. No changes were applied on the model, including not on its standard width.
ViT-B/16: In our implementation, we utilized the Vision Transformer (ViT) architecture, specifically the ViT-B/16 model, as described in [15]. The ViT-B/16 structure begins with an initial patch embedding layer, followed by a sequence of 12 transformer blocks. We used the ViT-B/16 implementation of PyTorch.
F.2 CKA Experimental Details
The CKA evaluations in this paper are based on two different datasets: CIFAR-10 and Tiny ImageNet.
  • CIFAR-10:
    • The CKA evaluations used a total of 2048 test examples, divided into 4 batches of 512 samples each.
    • Each batch contained almost an equal number of examples from each class.
    • The average CKA score was calculated over all batches to ensure robust and reliable evaluation.
    • In our experiments, we did not find a significant difference when compared to using a larger set of examples.
  • Tiny ImageNet:
    • The CKA evaluations used a total of 10,000 test examples.
    • No Label Noise: As described previously, the evaluations used 5 batches of 2000 samples each, i.e., a total of 10,000 test examples.
    • 20% Label Noise: For Tiny ImageNet with 20% label noise, the evaluation used a single batch of 2000 samples, i.e., a total of 2000 test examples.
The CKA evaluations considered the following epoch grids:
  • For ResNet-18 and ResNet-50 training on CIFAR-10, the epoch grid is always full with all the epochs from 0 to 3999.
  • For ResNet-18 training on Tiny ImageNet, the epoch grid includes all epochs from 0 to 300, every third epoch between 300 to 900, and every tenth epoch from 900 to 3999.
  • For ViT-B/16, the epoch grid includes all epochs from 0 to 300, every third epoch between 300 to 900, and every tenth epoch from 900 to 3999.
F.3 Decision Regions Similarity (DRS) Experimental Details
For evaluating the Decision Region Similarity (DRS) between two models, we used the following procedure:
  1. 1.
    First, we defined a grid of epochs from the 4000 training epochs where we sampled the DRS score.
    • For ResNet-18, the grid included all epochs from 0 to 300, every third epoch between 300 to 900, and every fifth epoch from 900 to 3999.
    • For ViT-B/16, the grid was the same, except for epochs 900-4000 where we sampled every tenth epoch.
  2. 2.
    For each epoch, we used the model from that epoch to extract the representations of the training set for the layer of interest. We then trained a linear classifier probe on top of the representations of the training set (all of the 50,000 train examples of CIFAR-10) and used the probe to predict the labels of the training set. All linear classifiers were trained with the Adam optimizer with a learning rate of 0.0001, cross-entropy loss, and for 10 epochs.
  3. 3.
    We randomly chose 500 triplets of training examples (these triplets were constant for all DRS experiments). Then, for each triplet, we computed the plane that they span (as mathematically explained in [12] and based on their code). The $j^{\text{th}}$ plane spanned by those 3 samples, denoted as $\mathcal{Q}_{j}\subset\mathcal{X}$, is a 2D plane in the input space. We then discretized each plane $\mathcal{Q}_{j}$ into a uniformly-spaced 2D grid of $|\mathcal{Q}_{j}^{\text{discrete}}|=2500$ points.
  4. 4.
    The DRS was then computed as the average agreement of the linear classifier probes on the discretized planes, according to Eq. (5).
Appendix G Summary of Experiment Settings
  • CKA experiments: Please refer to Table 1 for a summary of our CKA experiments, describing the datasets, architectures, optimizers, width parameters $k$, label noise percentage, and the corresponding figures of the results.
  • DRS experiments: As detailed in Section C, generating the Decision Region Similarity (DRS) heatmaps is highly computationally demanding, requiring significant time and compute resources. Therefore, we limited the DRS evaluations to the following configurations:
    • ResNet-18 trained on CIFAR-10 without label noise (Fig. 4).
    • ResNet-18 trained on CIFAR-10 with 20% label noise (Fig. 5).
    • ViT-B/16 trained on CIFAR-10 without label noise (Fig. 18).
Dataset Architecture Optimizer Width P. ($k$) % Label Noise Fig. #
CIFAR-10 ResNet-18 Adam 64 0 1
10 7
20 2
52 20 10
42 20 9
32 20 8
SGD+Mom. 64 20 14
SGD+Mom. 64 0 13
SGD (0.01) 64 20 15
Adam (wd) 64 0 16
ResNet-50 Adam N.A 20 12
ViT-B/16 Adam N.A 0 3
ViT-B/16 Adam N.A 20 17
Tiny ImageNet ResNet-18 Adam 64 0 11
20 20
Tiny ImageNet $1/2$ ResNet-18 Adam 64 20 21
Table 1: Summary of the CKA experiments conducted. Note that here Adam (wd) stands for training with Adam optimizer with weight decay regularization of 0.001, and Tiny ImageNet $1/2$ stands for the experminet where we used only half of the train samples of Tiny ImageNet. In addition, SGD (0.01) refers to Fig. 15, where we used SGD optimizer with constant learning rate of 0.01 and no momentum.
Appendix H Additional Experimental Results
Figure 7: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 10% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.
Figure 8: CKA evaluations for ResNet-18 of width parameter $k=32$ (which is narrower than the standard ResNet-18 width of $k=64$), trained on CIFAR-10 with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. (g) shows the test and train errors during training.
Figure 9: CKA evaluations for ResNet-18 of width parameter $k=42$ (which is narrower than the standard ResNet-18 width of $k=64$), trained on CIFAR-10 with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. (g) shows the test and train errors during training.
Figure 10: CKA evaluations for ResNet-18 of width parameter $k=52$ (which is narrower than the standard ResNet-18 width of $k=64$), trained on CIFAR-10 with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. (g) shows the test and train errors during training.
Figure 11: CKA evaluations for ResNet-18 (of standard width), trained on Tiny ImageNet (image size $32\times 32\times 3$) without label noise.
Figure 12: CKA evaluations for ResNet-50, trained on CIFAR-10 with 20% label noise.
Figure 13: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 without label noise. Here, the optimizer is SGD with constant learning rate 0.0001 and momentum of 0.9
Figure 14: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Here, the optimizer is SGD with constant learning rate 0.0001 and momentum of 0.9
Figure 15: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Here, the optimizer is SGD with constant learning rate 0.1 without momentum
Figure 16: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 with 20% label noise. Here, the optimizer is Adam with weight decay of 0.001.
Figure 17: CKA evaluations for ViT-B/16, trained on CIFAR-10 with 20% label noise.
Figure 18: Decision regions similarity (DRS) of linear classifier probes. Evaluations for ViT-B/16, trained on CIFAR-10 without label noise. Each subfigure shows the DRS of a specific layer in the ResNet-18 during its training. The DNN error curves are provided in Fig. 3g.
H.1 CKA Similarity of the First Layer Representations to Random Representations
As mentioned in Section 3.3 of the main text, the similarity of the first layer representations to the random representations (as appear, e.g., in Figs. 2a, 1a) is not only to the specific random initialization of the training process, but also to other random representations unrelated to the specific training process. This can be observed in Fig. 19 where we evaluate the CKA similarity of the first layer representations during training with respect to a random representation that was not used as the initialization of the specific training process; yet, it should be noted that the CKA similarity to the specific initialization is 1%-3% higher than the CKA similarity to another initialization.
Figure 19: First row from the CKA diagram in Fig. 2a for the first convolutional layer (blue line), together with the CKA similarity with respect to another random initialization (orange line). The model is ResNet-18 trained on CIFAR-10 with 20% label noise.
H.2 Additional CKA Evaluations
In Figs. 7-17 we provide CKA evaluations for various learning settings, in addition to the results in the main paper. See Table 1 for the summary of training settings.
H.3 Additional DRS Evaluations
In Fig. 18 we provide DRS evaluation for the ViT-B/16 model. This is in addition to the DRS evaluations on ResNet-18 in the main paper.
Figure 20: CKA evaluations for ResNet-18 (of standard width), trained on Tiny ImageNet with 20% label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.
Figure 21: CKA evaluations for ResNet-18 (of standard width), trained on Tiny-ImageNet with 20% label noise. In contrast to Fig. 20g, here we used only half of the train set (but still all the 200 classes) for training the model. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.